The price of a unit of capital is 2,000. The rate of depreciation is 1% per year. The annual real rate of interest is 14%.

What is the user cost of capital?

Please show working out
Answer #1

Price of one unit of capital = 2,000

Rate of depreciation = 1% or 0.01

Annual real rate of interest = 14% or 0.14

Calculate the user cost of capital -

User cost of capital = (rate of depreciation + annual real rate of interest) * price of one unit of capita

User cost of capital = (0.01 + 0.14) * 2,000

User cost of capital = 0.15 * 2,000 = 300

Thus,

The user cost of capital is 300.